# Direct and Converse Theorems in Estimating Signals with Sublinear Sparsity

### v1

@@ V Conclusions
Numerical results for AMP imply that AMP for sublinear sparsity has a performance bottleneck in another part—such as finite size effect—rather than the type of denoisers. To improve the convergence property of AMP, we need to deal with non-Gaussianity for estimation errors before denoising.
@@ Appendix A Estimators
A-A Proof of Proposition 1
We represent the $k$-sparse vector $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$ with the support set $\mathcal{S}_{\boldsymbol{x}}\in\mathcal{X}_{k}^{N}(\{1\})$ and the non-zero signals $\boldsymbol{u}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}\in\mathcal{U}^{k}$. Using the definition of $\boldsymbol{y}$ in (1), we write the ML estimator (6) as
@@ A-A Proof of Proposition 1
Let $\boldsymbol{u}_{+}$ and $\boldsymbol{u}_{-}$ denote the vectors obtained by eliminating the negative and positive elements from $\boldsymbol{u}$, respectively. For the dimension $k_{0}\in\{0,\ldots,k\}$ of $\boldsymbol{u}_{+}$, we decompose $\mathcal{U}^{k}$ into $k+1$ disjoint subsets $\mathcal{U}^{k}(k_{0})=\{\boldsymbol{u}\in\mathcal{U}^{k}:\boldsymbol{u}_{+}\in\mathcal{U}_{+}^{k_{0}}\}$ with $\mathcal{U}_{+}=\{u_{m}\in\mathcal{U}:u_{m}>0\}$ and $\mathcal{U}_{-}=\{u_{m}\in\mathcal{U}:u_{m}<0\}$. We represent the objective function in the minimization problem (37) as
@@
A-B Non-Separable Bayesian Estimator
We derive the non-separable Bayesian estimator (34). We start with the optimum posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$, given by
$\hat{x}_{\mathrm{opt},n}=\frac{\mathbb{E}_{\boldsymbol{x}}[x_{n}p(\boldsymbol{y}|\boldsymbol{x})]}{\mathbb{E}_{\boldsymbol{x}}[p(\boldsymbol{y}|\boldsymbol{x})]}.$ (43)
Using the definition of the AWGN channel (1) yields
@@ A-B Non-Separable Bayesian Estimator
$=\frac{|\mathcal{M}|^{-1}\sum_{m\in\mathcal{M}}u_{m}e^{-\frac{(y_{n}-u_{m})^{2}}{2\sigma_{N/k}^{2}}}}{|\mathcal{M}|^{-1}\sum_{m\in\mathcal{M}}e^{-\frac{(y_{n}-u_{m})^{2}}{2\sigma_{N/k}^{2}}}+e^{-\frac{y_{n}^{2}}{2\sigma_{N/k}^{2}}}e^{L_{\mathrm{opt}}(\boldsymbol{y}_{\setminus n})}},$ (44)
@@ B-A Overview
The proof of Theorem 1 consists of three steps. A first step is the decomposition of the type $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ in (12) into two types: A first type is associated with the true signal support $\mathcal{S}_{\boldsymbol{x}}$. The other type corresponds to the remaining indices $\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}$. For the former type, the simple union bound is used.
@@ B-B Decomposition
For a subset $\mathcal{S}\subset\mathcal{N}$ of size $k$, we decompose a $k$-sparse vector $\boldsymbol{x}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S})$ in (12) into $\boldsymbol{x}_{\mathcal{S}}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ and $\boldsymbol{x}_{\mathcal{N}\setminus\mathcal{S}}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ for $\boldsymbol{m}_{1}\in\mathcal{M}^{k-w}$ and $\boldsymbol{m}_{2}\in\mathcal{M}^{w}$ satisfying $\boldsymbol{m}=(\boldsymbol{m}_{1},\boldsymbol{m}_{2})$, with
@@
Fix $w\in\{0,\ldots,k\}$, $w^{\prime}\in\{0,\ldots,k-w\}$, $\boldsymbol{m}\in\mathcal{M}^{k}$, and $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$. If $\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ and $N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}$ hold, there is some $\boldsymbol{x}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ such that $p(\boldsymbol{y}|\boldsymbol{x}^{\prime})\geq p(\boldsymbol{y}|\boldsymbol{x})$ and $N_{\boldsymbol{x},\boldsymbol{x}^{\prime}}=w^{\prime}$ are satisfied. Decompose $\boldsymbol{x}^{\prime}$ into the two vectors $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{x}^{\prime}_{\mathcal{S}_{\boldsymbol{x}}}$ and $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{x}^{\prime}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}$. Since $N_{\boldsymbol{x},\boldsymbol{x}^{\prime}}$ is a deterministic function of $\boldsymbol{x}_{1}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}$ and $\boldsymbol{x}_{1}^{\prime}$, we rewrite $N_{\boldsymbol{x},\boldsymbol{x}^{\prime}}$ as $N_{\boldsymbol{x}_{1},\boldsymbol{x}_{1}^{\prime}}$. Using the union bound with respect to $\boldsymbol{x}_{1}^{\prime}$ yields
@@ B-C Gallager’s Bound
To evaluate the upper bound (53), we focus on each term in the summation. Using the definition of the AWGN channel (1) yields
@@ D-B KL Divergence
Let
@@
From the definition of $\boldsymbol{y}$ in (1), the true expectation $\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]$ is equal to $\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]+N\sigma_{N/k}^{2}=N\lambda\sigma_{N/k}^{2}$. By definition, $q(\boldsymbol{y})$ is represented as
@@
We note that $u$ satisfying $|u|=u_{\mathrm{max}}$ is the maximizer of $h(u,\omega)$ over $u\in\mathcal{U}\cup\{0\}$. Thus, we have the following uniform upper bound $G_{k-w}^{k}(\boldsymbol{x}_{1},\boldsymbol{\omega}_{1})\leq G_{k-w}^{k}(\boldsymbol{\omega}_{1})$, with $G_{0}^{k}(\boldsymbol{\omega}_{1})=1$ and
@@ D-D Non-Typical Noise Samples
It is difficult to evaluate the expectation in (122) with respect to $\boldsymbol{\omega}_{1}$. Furthermore, using Jensen’s inequality results in a loose upper bound on $g_{2,i}$. To circumvent this difficulty, we prove the following upper bound on $G_{k-w}^{k}(\boldsymbol{\omega}_{1})$ in (98) conditioned on $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}$:
@@ D-E Second Term in (127)
We next focus on $\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\})$. Decomposing the event $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}$ yields
@@ D-F First Term in (127)
$g_{2,i,j}^{(1)}=\mathbb{E}\left[1((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\})\log\left\{|\mathcal{M}|^{-(k-j)}G_{k-j}^{k-j}\right\}\right].$ (144)
Decomposing $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}$ into two disjoint events $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\in\mathcal{I}\}$ and $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\notin\mathcal{I}\}$, we represent $g_{2,i,j}^{(1)}$ with $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}$ replaced by the former and latter events as $g_{2,i,j}^{(1,1)}$ and $g_{2,i,j}^{(1,2)}$, respectively. By definition, we have $g_{2,i,j}^{(1)}=g_{2,i,j}^{(1,1)}+g_{2,i,j}^{(1,2)}$. The goal is evaluation of $g_{2}^{(1)}=g_{2}^{(1,1)}+g_{2}^{(1,2)}$ with $g_{2}^{(1,l)}=\sum_{i=1}^{k}g_{2,i}^{(1,l)}$ and $g_{2,i}^{(1,l)}=\sum_{j=0}^{k}g_{2,i,j}^{(1,l)}$ for $l\in\{1,2\}$.

### v3

@@ V Conclusions
Numerical results for AMP suggest that AMP for sublinear sparsity has a performance bottleneck in another part—such as finite size effect—rather than the type of denoisers. To improve the convergence property of AMP, we need to deal with non-Gaussianity for estimation errors before denoising.
@@ Appendix A Estimators
A-A Separable Bayesian Estimator
[…]
Let $U\in\mathbb{R}\setminus\{0\}$ denote a random variable representing i.i.d. non-zero signals. As derived in [1, Section IV-A], the element-wise posterior mean estimator $\mathbb{E}[x_{n}|y_{n}]$ is given by
$\mathbb{E}[x_{n}|y_{n}]=\frac{k\mathbb{E}_{U}[e^{-(y_{n}-U)^{2}/(2\sigma_{N/k}^{2})}]\mathbb{E}[U|U+\omega_{n}=y_{n}]}{k\mathbb{E}_{U}[e^{-(y_{n}-U)^{2}/(2\sigma_{N/k}^{2})}]+(N-k)e^{-y_{n}^{2}/(2\sigma_{N/k}^{2})}},$ (39)
with $\omega_{n}\sim\mathcal{N}(0,\sigma_{N/k}^{2})$ independent of $U$. For the constant non-zero signals $\mathbb{P}(U=1)=1$, the element-wise posterior mean estimator reduces to
$\mathbb{E}[x_{n}|y_{n}]=\frac{ke^{-(y_{n}-1)^{2}/(2\sigma_{N/k}^{2})}}{ke^{-(y_{n}-1)^{2}/(2\sigma_{N/k}^{2})}+(N-k)e^{-y_{n}^{2}/(2\sigma_{N/k}^{2})}}.$ (40)
To compare the main results in this paper with an existing result [1, Lemma 1] for the separable Bayesian estimator, we focus on discrete non-zero signals $\mathcal{U}=\{u_{m}\in\mathbb{R}\setminus\{0\}:m\in\mathcal{M}\}$. We do not necessarily assume the uniform distribution on $\mathcal{U}$. The separable Bayesian estimator (39) was proved to achieve the following square error in the sublinear sparsity limit:
Theorem 3 ([1])
For all $\sigma^{2}>0$, we have
$\frac{1}{k}\sum_{n\in\mathcal{N}}\mathbb{E}\left[\left(x_{n}-\mathbb{E}[x_{n}|y_{n}]\right)^{2}\right]\to\mathbb{E}\left[U^{2}1(U^{2}<2\sigma^{2})\right]$ (41)
in the sublinear sparsity limit.
Proof:
Theorem 3 is equivalent to [1, Lemma 1]. More precisely, [1, Lemma 1] claimed the convergence in probability to treat random small perturbation. When the random small perturbation is eliminated, however, the convergence in mean is what was actually proved in [1].
To use [1, Lemma 1], we need to confirm Assumption 7 in [1, p. 4608]. For the discrete random variable $U\in\mathcal{U}$, Assumption 7 is trivial with the exception of the uniform Lipschitz-continuity for $\mathbb{E}[U|U+\omega_{n}=y_{n}]$ with respect to $\sigma_{N/k}^{2}>0$. In fact, this uniform Lipschitz-continuity is broken for discrete $U\in\mathcal{U}$ as $\sigma_{N/k}^{2}\to 0$. Nonetheless, we can still use [1, Lemma 1]: The uniform Lipschitz-continuity assumption was used in [1] only to justify that there are some constants $L>0$ and $C>0$ such that $|\mathbb{E}[U|U+\omega_{n}=y_{n}]|\leq L|y_{n}|+C$ is satisfied for all $y_{n}\in\mathbb{R}$. Since $\mathbb{E}[U|U+\omega_{n}=y_{n}]$ is uniformly bounded for discrete $U\in\mathcal{U}$, the inequality to be justified is trivial for discrete $U\in\mathcal{U}$. ∎
Theorem 3 implies that the separable Bayesian estimator achieves vanishing square error for all $\sigma^{2}<u_{\mathrm{min}}^{2}/2$, as proved in Theorem 1 for the ML estimator. For all $\sigma^{2}>u_{\mathrm{max}}^{2}/2$, on the other hand, the square error is equal to $\mathbb{E}[U^{2}]$. The latter result is consistent to Theorem 2. In the intermediate region $2\sigma^{2}\in(u_{\mathrm{min}}^{2},u_{\mathrm{max}}^{2})$, Theorem 3 indicates that the square error is in the open interval $(0,\mathbb{E}[U^{2}])$.
A-B Proof of Proposition 1
We represent the $k$-sparse vector $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$ with the support set $\mathcal{S}_{\boldsymbol{x}}\in\mathcal{X}_{k}^{N}(\{1\})$ and the non-zero values $\boldsymbol{u}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}\in\mathcal{U}^{k}$. Using the definition of $\boldsymbol{y}$ in (1), we write the ML estimator (6) as
@@ A-B Proof of Proposition 1
Let $\boldsymbol{u}_{+}$ and $\boldsymbol{u}_{-}$ denote the vectors obtained by extracting the positive and negative elements from $\boldsymbol{u}$, respectively. For the dimension $k_{0}\in\{0,\ldots,k\}$ of $\boldsymbol{u}_{+}$, we decompose $\mathcal{U}^{k}$ into $k+1$ disjoint subsets $\mathcal{U}^{k}(k_{0})=\{\boldsymbol{u}\in\mathcal{U}^{k}:\boldsymbol{u}_{+}\in\mathcal{U}_{+}^{k_{0}}\}$ with $\mathcal{U}_{+}=\{u_{m}\in\mathcal{U}:u_{m}>0\}$ and $\mathcal{U}_{-}=\{u_{m}\in\mathcal{U}:u_{m}<0\}$. We represent the objective function in the minimization problem (42) as
@@
A-C Non-Separable Approximate Bayesian Estimator
We present a heuristic derivation for the non-separable approximate Bayesian estimator (36). We start with the optimum posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$, given by
$\hat{x}_{\mathrm{opt},n}=\frac{\mathbb{E}_{\boldsymbol{x}}[x_{n}p(\boldsymbol{y}|\boldsymbol{x})]}{\mathbb{E}_{\boldsymbol{x}}[p(\boldsymbol{y}|\boldsymbol{x})]},$ (48)
where the expectation is over $\boldsymbol{x}$ uniformly distributed on $\mathcal{X}_{k}^{N}(\mathcal{U})$. Using the definition of the AWGN channel (1) yields
@@ A-C Non-Separable Approximate Bayesian Estimator
$=\frac{\sum_{m\in\mathcal{M}}u_{m}e^{-\frac{(y_{n}-u_{m})^{2}}{2\sigma_{N/k}^{2}}}}{\sum_{m\in\mathcal{M}}e^{-\frac{(y_{n}-u_{m})^{2}}{2\sigma_{N/k}^{2}}}+e^{-\frac{y_{n}^{2}}{2\sigma_{N/k}^{2}}}e^{L_{\mathrm{opt}}(\boldsymbol{y}_{\setminus n})}},$ (49)
@@ B-A Overview
The proof of Lemma 1 consists of three steps. A first step is the decomposition of the type $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ in (12) into two types: A first type is associated with the true signal support $\mathcal{S}_{\boldsymbol{x}}$. The other type corresponds to the remaining indices $\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}$. For the former type, the simple union bound is used.
@@ B-B Decomposition
For a subset $\mathcal{S}\in\mathfrak{S}_{k}^{N}$ of size $k$, we decompose a $k$-sparse vector $\boldsymbol{x}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S})$ in (12) into $\boldsymbol{x}_{\mathcal{S}}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ and $\boldsymbol{x}_{\mathcal{N}\setminus\mathcal{S}}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ for $\boldsymbol{m}_{1}\in\mathcal{M}^{k-w}$ and $\boldsymbol{m}_{2}\in\mathcal{M}^{w}$ satisfying $\boldsymbol{m}=(\boldsymbol{m}_{1},\boldsymbol{m}_{2})$, with
@@
Fix $w\in\{0,\ldots,k\}$, $w^{\prime}\in\{0,\ldots,k-w\}$, $\boldsymbol{m}\in\mathcal{M}^{k}$, and $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$. If $\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ and $N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}$ hold, there is some $\boldsymbol{x}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ such that $p(\boldsymbol{y}|\boldsymbol{x}^{\prime})\geq p(\boldsymbol{y}|\boldsymbol{x})$ and $N_{\boldsymbol{x},\boldsymbol{x}^{\prime}}=w^{\prime}$ are satisfied. Decompose $\boldsymbol{x}^{\prime}$ into the two vectors $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{x}^{\prime}_{\mathcal{S}_{\boldsymbol{x}}}$ and $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{x}^{\prime}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}$. Since $N_{\boldsymbol{x},\boldsymbol{x}^{\prime}}$ is a deterministic function of $\boldsymbol{x}_{1}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}$ and $\boldsymbol{x}_{1}^{\prime}$, we rewrite $N_{\boldsymbol{x},\boldsymbol{x}^{\prime}}$ as $N_{\boldsymbol{x}_{1},\boldsymbol{x}_{1}^{\prime}}$.
We evaluate $\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$, in which the distribution is induced from that of $\boldsymbol{y}$ conditioned on $\boldsymbol{x}$ in (1). Using the union bound with respect to $\boldsymbol{x}_{1}^{\prime}$ yields
@@ B-C Gallager’s Bound
To evaluate the upper bound (58), we focus on each term in the summation. Using the definition of the AWGN channel (1), as evaluated in [30, Chapter 7], we obtain
@@ D-B KL Divergence
We recall that the true marginal pdf of $\boldsymbol{y}$ is induced from the conditional pdf $p(\boldsymbol{y}|\boldsymbol{x})$ defined via (1) and the true distribution of $\boldsymbol{x}$—uniform distribution on $\mathcal{X}_{k}^{N}(\mathcal{U})$. Let
@@
From the definition of $\boldsymbol{y}$ in (1) and the independence between $\boldsymbol{x}$ and $\boldsymbol{y}$, the true expectation $\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]$ is equal to $\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]+N\sigma_{N/k}^{2}=N\lambda\sigma_{N/k}^{2}$. By definition, $q(\boldsymbol{y})$ is represented as
@@
We note that $u$ satisfying $|u|=u_{\mathrm{max}}$ is the maximizer of $h(u,\omega)$ over $u\in\mathcal{U}\cup\{0\}$. Thus, we have the uniform upper bound $G_{k-w}^{k}(\boldsymbol{x}_{1},\boldsymbol{\omega}_{1})\leq G_{k-w}^{k}(\boldsymbol{\omega}_{1})$, with $G_{0}^{k}(\boldsymbol{\omega}_{1})=1$ and
@@ D-D Non-Typical Noise Samples
It is difficult to evaluate the expectation in (127) with respect to $\boldsymbol{\omega}_{1}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{k})$. Furthermore, using Jensen’s inequality results in a loose upper bound on $g_{2,i}$. To circumvent this difficulty, we prove the following upper bound on $G_{k-w}^{k}(\boldsymbol{\omega}_{1})$ in (103) conditioned on $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}$:
@@ D-E Second Term in (132)
We next focus on $\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\})$, in which the probability distribution is induced from $\boldsymbol{\omega}_{1}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{k})$ through $\mathcal{I}$ in (128) and $\mathcal{E}_{i}^{k-1}=\{\Omega_{i}^{k-1}:|\Omega_{i}^{k-1}|\leq d_{i}^{k-1}\}$ with $\Omega_{i}^{k-1}=(\boldsymbol{x}_{i}^{k-1})^{\mathrm{T}}\boldsymbol{\omega}_{1}/\sigma_{N/k}^{2}$. Decomposing the event $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}$ yields
@@ D-F First Term in (132)
$g_{2,i,j}^{(1)}=\mathbb{E}\left[1((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\})\log\left\{|\mathcal{M}|^{-(k-j)}G_{k-j}^{k-j}\right\}\right],$ (149)
where the probability distribution is induced from $\boldsymbol{\omega}_{1}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{k})$ through $\mathcal{I}$ in (128) and $\mathcal{E}_{i}^{k-1}=\{\Omega_{i}^{k-1}:|\Omega_{i}^{k-1}|\leq d_{i}^{k-1}\}$ with $\Omega_{i}^{k-1}=(\boldsymbol{x}_{i}^{k-1})^{\mathrm{T}}\boldsymbol{\omega}_{1}/\sigma_{N/k}^{2}$.
We decompose $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}$ into two disjoint events $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\in\mathcal{I}\}$ and $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\notin\mathcal{I}\}$. Representing $g_{2,i,j}^{(1)}$ with $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}$ replaced by the former and latter events as $g_{2,i,j}^{(1,1)}$ and $g_{2,i,j}^{(1,2)}$, respectively, we have $g_{2,i,j}^{(1)}=g_{2,i,j}^{(1,1)}+g_{2,i,j}^{(1,2)}$. The goal is evaluation of $g_{2}^{(1)}=g_{2}^{(1,1)}+g_{2}^{(1,2)}$ with $g_{2}^{(1,l)}=\sum_{i=1}^{k}g_{2,i}^{(1,l)}$ and $g_{2,i}^{(1,l)}=\sum_{j=0}^{k}g_{2,i,j}^{(1,l)}$ for $l\in\{1,2\}$.